### v1

# Determination of galaxy photometric redshifts using Conditional Generative Adversarial Networks (CGANs)

Abstract
Accurate and reliable photometric redshifts determination is one of the key aspects for wide-field photometric surveys. Determination of photometric redshift for galaxies, has been traditionally solved by use of machine-learning and artificial intelligence techniques trained on a calibration sample of galaxies, where both photometry and spectrometry are determined. On this paper, we present a new algorithmic approach for determining photometric redshifts of galaxies using Conditional Generative Adversarial Networks (CGANs). Proposed CGAN implementation, approaches photometric redshift determination as a probabilistic regression, where instead of determining a single value for the estimated redshift of the galaxy, a full probability density is computed. The methodology proposed, is tested with data from Dark Energy Survey (DES) Y1 data and compared with other existing algorithm such as a Random Forest regressor.
keywords:
Conditional Generative Adversarial Networks , photometric-redshift , galaxy-surveys
journal: Astronomy and Computing
[…]
[mgf]organization=School of Architecture, Engineering and Design, Universidad Europea de Madrid,addressline=Calle Tajo s/n, city=Villaviciosa de Odon, postcode=28670, state=Madrid, country=Spain
1 Introduction
Wide-field photometric surveys have been a major source for experimental results for Observational Cosmology. Among many of the present and future photometric surveys we can find: DES [^2] https://www.darkenergysurvey.org, LSST [^3] https://www.lsst.org, PAU [^4] https://pausurvey.org, J-PAS [^5] https://www.j-pas.org and Euclid [^6] https://www.euclid-ec.org. One of the key aspects of wide-fied photometric surveys is the reliable determination of the redshift of the galaxies. At photometric surveys, the redshift of galaxy spectra is inferred by measuring the brightness of galaxies at broad-band filters instead of determining the Doppler shift of their spectra with a high-resolution spectrometer.
The usual approach for translating from the brightness measured at the broad-band filters to a redshift that is the closest as possible to the potential actual spectroscopic redshift, the use of machine-learning and artificial intelligence techniques are used. These techniques, make use of a calibration sample of galaxies with known both the photometry and the high-resolution spectra. This calibration sample, is used by the machine-learning algorithms as a source for indenfying and discovering patterns relating the brightness at the different broad-band filters with the spectroscopic redshift.
Previous machine-learning algorithms for photometric redshift determination have included: template-fitting (Schaap and van de Weygaert, 2000; Bolzonella et al., 2011; Benítez, 2011; Arnouts and Ilbert, 2011; Laur et al., 2022), neural networks (Collister and Lahav, 2004; Sadeh et al., 2016; Mahmud Pathi et al., 2024; Hoyle, 2016; Chunduri and Mahesh, 2023; Carrasco Kind and Brunner, 2013; Almosallam et al., 2016; Cavuoti et al., 2017), boosted decision trees (Gerdes et al., 2010)-, convolutional neural networks (D’Isanto and Polsterer, 2018; Schuldt et al., 2021), bayesian neural networks (Lima et al., 2022), random forest (Lu et al., 2023), recurrent neural networs (Luo et al., 2024) and nearest neighbours (Graham et al., 2018). A good systematic review on the different types of photometric redshift algorithms can be found at Sánchez et al. (2014); Salvato et al. (2019) and Newman and Gruen (2022).
As it can be identified at literature review, the most widely used machine-learning technique at photometric redshift estimation are neural networks at their many flavours (dense, recurrent and convolutional), with a strong presence of feed-forward neural networks composed by dense layers. Feed-forward neural networks composed by dense layers are a great computational technique for solving the photometric redshift estimation task, as they constitute an excellent regression technique for multivariate analysis that -if provided a proper loss function-, can minimize the errors between the inferred redshift and the actual spectroscopic redshift of the calibration sample without the need for providing an explicit functional dependence between, template or prior assumption (Bishop, 2006; Goodfellow et al., 2016). Nevertheless, the estimation of these photometric redshifts constitute a point estimate of the spectroscopic redshift instead of a probability density function. A good point estimate is necessary for some cosmological probes -such as Baryon Acoustic Oscillation (Crocce et al., 2018)-, but for other cosmological probes require to have a good control of the tails of the probability distribution function -such as number-counts magnification (Garcia-Fernandez et al., 2018)-.
A new class of neural network that has still not been explored at photometric redshift estimation are the Generative Adversarial Networks (GANs). GANs (Goodfellow et al., 2014), are a set of two neural networks, a Generator network and a Discriminator network that compete between each other. The Generator aims to generate synthetic data mimicking the real data, whereas the Discriminator network aims to identify if a given data-record is synthetic data produced by the Generator or is an observation of the real data. On the process of training this algorithm, both Generator and Discriminator compete with each other in a zero-sum competitive game approach, that is finished when the Generator can completely fool the Discriminator, so the Discriminator is not able to properly disentangle synthetic data from real data. As a consequence of this, the Generator network can accurately map the full probability distribution of the underlying data without the need of providing any prior assumption or template.
A very relevant special case for GANs, are the Conditional Generative Adversarial Networks (CGANs), which instead of tracing the full probability density function of the underlying data, they trace the conditional probability density function, provided some input condition (Mirza and Osindero, 2014).
On this paper, we propose the use of CGANs for estimating photometric redshifts using the magnitudes measured at broad-band filters. Proposed algorithm is tested with data from the Dark Energy Survey Y1 data that is overlapping with SDSS Stripe-82. Results obtained by proposed CGAN are compared with a classic machine-learning technique such as a Random Forest regressor.
2 Methodology
Let ${\bf x}_{i}$ be a sample of photometric data from magnitudes measured at broad-band pass filters for the $i$-th galaxy of some data and $y_{i}$ its corresponding known spectroscopic redshift, such that $\{({\bf x}_{i},y_{i})\}_{i=1}^{N_{train}}$ constitutes the data training set. Let ${\bf z}_{i}$ be a vector of randomly-generated numbers [^7] Beware that Computer Science standard notation at literature on GANs, name the random vectors with $z$. Do not confuse this $z$ with the redshift of the galaxies (that will be either denoted in this paper as $y$ or $\hat{y}$ for spectroscopic or photometric redshifts respectively., associated to the $i-th$ galaxy.
The generator network $G$ constitutes a function such that $\hat{y}_{i}=G({\bf z}_{i}|{\bf x}_{i};\theta_{G})$, where $\hat{y}_{i}$ is the estimated photometric redshift for the set of magnitudes ${\bf x}_{i}$ and $\theta_{G}$ are the weights defining the generator neural network. On the other side, the discriminator network $D$ provides a function such that $d_{i}=D(\hat{y}_{i};\theta_{D})$with $d_{i}\in[0,1]$- is a classifier that identifies if $\hat{y}_{i}$ is a real data from the training sample or a synthetic-generated data produced by the generator network and $\theta_{D}$ are the weights defining the discriminator network.
The process of training a GAN network constitutes a min-max problem such that $\min_{\theta_{G}}\max_{\theta_{D}}V(D,G)$. The choice of the function $V(D,G)$ is a vast problem within the field of Computer Science. As demonstrated by Nowozin et al. (2016), any GAN can can be interpreted as a special type of variational divergence estimation. Thus, the function $V(G,D)$ can be placed on the most generic formulation as
$V(D,G)=\mathbb{E}_{{\bf x}\sim p_{d}({\bf x})}[\mathbb{E}_{y\sim p_{d}(y)}[g_{f}(D(y|{\bf x}))]+\mathbb{E}_{{\bf z}\sim p_{z}({\bf z})}[-f^{*}(g_{f}(D(G({\bf z}|{\bf x}))))]],$ (1)
where $g_{f}$ denotes the output activation function and $f^{*}$ is the corresponding Fenchel conjugate function of $g_{f}$ (Hiriart-Urruty and Lemaréchal, 2001). The functions $g_{f}$ and $f^{*}$ can be chosen freely, provided they are derived from any $f$-divergence (Csiszár and Shields, 2004; Liese and Vajda, 2006; Nguyen et al., 2007; Reid and Williamson, 2011). Conversely, $\mathbb{E}_{{\bf x}\sim p_{d}({\bf x})}$, $\mathbb{E}_{y\sim p_{d}(y)}$ and $\mathbb{E}_{{\bf z}\sim p_{d}({\bf z})}$ denote the expected values over $\bf x$, $y$ and $\bf z$ respectively.
Taking this into account, the loss function to be minimized for the discriminator network is given by
$\mathcal{L}_{D}(\theta_{D})=\frac{-1}{N_{batch}}\sum\limits_{i=1}^{N_{batch}}[g_{f}(D(y_{i}|{\bf x}_{i};\theta_{D})-f^{*}(g_{f}(D(G({\bf z}_{i}|{\bf x}_{i});\theta_{G})|{\bf x}_{i};\theta_{D})],$ (2)
whereas the loss for the generator to be minimized simultaneously is given by
$\mathcal{L}_{G}(\theta_{G})=\frac{-1}{N_{batch}}\sum\limits_{i=1}^{N_{batch}}g_{f}(D(G({\bf z}_{i}|{\bf x}_{i};\theta_{G})|{\bf x}_{i};\theta_{D}).$ (3)
From all the possible $f$-divergences, we select the Kullback-Leibler divergence (Kullback and Leibler, 1951), given by
[…]
Thus, by using the KL-divergence as the proposed $f$-divergence, the corresponding $g_{f}(x)$ and $g(f^{*}(x))$ are given by Nowozin et al. (2016)
$g_{f}(x)=x$ (5)
and
$f^{*}(g_{f}(x))=e^{x-1}.$ (6)
By using this $f$-divergence approach, the photometric redshift inferred by the generator network -given a fixed set of values for the magnitudes ${\bf x}_{i}$-, will be a function of the random vector ${\bf z}_{j}$ such that
$\hat{y}_{i}({\bf z}_{j})=G({\bf z}_{j}|{\bf x}_{i}).$ (7)
If $N_{r}$ different random vectors are considered, the set of values $\{\hat{y}_{i}({\bf z}_{j})\}_{j=1}^{N_{r}}$, will be a random variable distributed according to the underlying conditional probability density of the spectroscopic redshift for the value of magnitudes ${\bf x}_{i}$. Thus, provided that $N_{r}$ is large, the distribution of $\hat{y}_{i}({\bf z_{j}})$ is an estimator of the probability density function of the true spectroscopic redshift for that single $i$-th galaxy.
Proposed topology of the neural network for the generator is a sequence of 3 fully connected layers, where the first two layers are followed by a Batch Normalization layer and a ReLU activation function. Input neurons of the first fully connected layer has $4+Z_{DIM}$ neurons with $G_{DIM}$ output neurons. Hidden fully connected layer consist of $G_{DIM}$ input neurons and $G_{DIM}$ output neurons. Final fully connected layer consist of $G_{DIM}$ input neurons and a single output neuron. The input of the generator network is composed by the vector $\bf x$ of the four galaxy magnitudes and the random vector $\bf z$. The output of the generator network is directly the inferred photometric redshift $\hat{y}$.
For the discriminator neural network a sequence of 3 fully connected layers, where the first two layers are followed by a Batch Normalization layer and a ReLU activation function. Input neurons of the first fully connected layer has $5$ neurons with $D_{DIM}$ output neurons. Hidden fully connected layer consist of $D_{DIM}$ input neurons and $D_{DIM}$ output neurons. Final fully connected layer consist of $D_{DIM}$ input neurons and a single output neuron followed by a sigmoid activation function, $\sigma(x)=1/(1+e^{-x})$. The input of the discriminator network is composed by a redshift (spectroscopic -$y$- or photometric -$\hat{y}$-), and the vector $\bf x$ of galaxy magnitudes. The output of the discriminator network is a decimal number between 0 and 1 that indicates the probability of the input galaxy to be real.
Topology of generator and discriminator neural networks can be seen at Figure 1.
Figure 1: Topology of the neural networks used. Blue boxes denote the fully connected layers, green boxes denote the normalization layers and red boxes denote the activation functions.
In order to compare the results of the proposed CGAN with classic machine-learning approaches, a Random Forest regressor is also employed (Breiman, 2001)
3 Data Analysis
Proposed CGAN was tested at DES-Y1 data spectroscopy-matched with SDSS galaxies. Data was taken from the public data releases DR1 (Abbott et al., 2018; Drlica-Wagner et al., 2018) from The Dark Energy Survey Collaboration that are available at the NCSA repository [^8] https://des.ncsa.illinois.edu/releases/y1a1.
Code for the CGAN was implemented under Python [^9] https://www.python.org using PyTorch [^10] https://pytorch.org (Paszke et al., 2019), whereas code for the Random Forest Regressor was implemented using Scikit-Learn [^11] https://scikit-learn.org (Pedregosa et al., 2011). Full code of this analysis can be found at author’s GitHub repository [^12] https://github.com/mgarciafernandez-uem/CGAN-photoz.
From the DES-Y1 proposed sample, we take the Stripe-82 subset of galaxies with matched spectroscopic redshift from SDSS. For consideration of the sample, we restrict ourselves to the galaxies with spectroscopic redshift $0.80.0<z_{sp}<$ 0.8. This quality cut is imposed to avoid a very long tail up to redshift 2, but with a few galaxies, as the inclusion of these elements on the training sample of the algorithm can potentially introduce biases due to the neural network infer magnitude-redshift relationships from an under-represented sample of galaxies of high-redshift. Selected measurements of magnitudes for the galaxies are the MAG_AUTO magnitudes at the $griz$ band-pass filters as measured by SExtractor (Bertin and Arnouts, 1996). The distribution of the $griz$ magnitudes and spectroscopic redshift of the selected calibration sample of galaxies can be seen along with their correlation plots at Figure 2. Total number of galaxies at the final calibration sample is 33410. From this final calibration sample, we do a split between a training sample and a test sample with 80% and 20% of the galaxies respectively. The training sample is used for determining the parameters of the photometric redshift algorithm during the training process. On the other side, the test sample is used for measuring the properties of the algorithm on a set of galaxies unseen by the photometric redshift algorithm.
Figure 2: Distribution and correlation of the $griz$ MAG_AUTO magnitudes and the measured spectroscopic redshift
Hyperparameters of the model defining the sizes of the dense layers of generator ($G_{DIM}$) and the discriminator ($D_{DIM}$) along with the size the random vector ($Z_{DIM}$) can be seen at Table 1. Training strategy used the Adam optimizer with a learning rate parameter of $lr=10^{-4}$ for both generator and discriminator network using $10000$ training epochs, using a step learning-rate approach reducing the learning rate by a factor of $0.2$ every $2000$ training epochs. Training strategy is the same for the generator and the discriminator networks. The loss for both generator and discriminator as a function of the training epoch can be seen at Figure 3, showing a good and stable convergence of the losses for both the generator and discriminator networks.
Hyperparameter Value
$Z_{DIM}$ 20
$D_{DIM}$ 32
$G_{DIM}$ 32
Table 1: Hyperparameters of the CGAN used for this paper.
Figure 3: Loss function for the generator and discriminator networks as a function of the training epoch as measured at the training and test data samples. Values of the measured losses have been shifted +1 in order to avoid negative values at the Y-axis. A plot showing the Mean Square Error (MSE) between the inferred photometric redshift and the true spectroscopic redshift is also displayed.
A comparison of the observed distribution of inferred redshift of galaxies for the test sample is compared with the true spectroscopic redshift can be seen at Figure 4. In addition, a comparison of the dispersion between the true spectroscopic redshift and the inferred photometric redshift can be see at Figure 5. From these plots, we can see that the photometric redshifts inferred by the proposed CGAN approach can accurately trace the true spectroscopic redshift distribution of the test sample of galaxies, reaching an agreement similar to the classic Random Forest Regressor. In addition, we can see that there is a similar level of dispersion on the photometric redshifts inferred by the CGAN approach and the Random Forest Regressor. It is important to remark that no significant biases or asymmetries are spotted. We see that dispersion of inferred photometric redshift randomly fluctuate around the true spectroscopic redshift for all the range of values of redshift.
Figure 4: Distribution of the photometric redshift of galaxies inferred from both proposed CGAN approach and the classic Random Forest algorithm. Both distributions are compared with the true spectroscopic redshift (black solid-line). Measurements are made on the test sample of galaxies.
Figure 5: Dispersion between the photometric redshift inferred by proposed CGAN and classic Random Forest algorithm and the true spectroscopic redshift. The black-dashed line displays those redshifts where the photometric redshift is equal to the spectroscopic redshift. Measurements are made on the test sample of galaxies.
Quality metrics for photometric redshift determination are based on the ones proposed at Sánchez et al. (2014) and can be found at Table 2. Used quality metrics for photometric comparison, are: the mean bias between the photometric redshift and the spectroscopic redshift ($\bar{\Delta z}$), the median of the bias ($\Delta z_{50}$), the standard deviation of the bias ($\sigma_{\Delta z}$), the fraction of galaxies with spectroscopic redshift outside of the two-standard deviation range ($out_{2\sigma}$) and the fraction of galaxies with spectroscopic redshift outside of the three-standard deviation range ($out_{3\sigma}$).
Parameter Description CGAN Random Forest
$\bar{\Delta z}$ Mean bias -0.0046 -0.0030
$\Delta z_{50}$ Median of the bias -0.0046 -0.0055
$\sigma_{\Delta z}$ Standard deviation of the bias 0.092 0.056
$out_{2\sigma}$ Outliers beyond 2 standard deviations 4% 5%
$out_{3\sigma}$ Outliers beyond 3 standard deviations 2% 2%
Table 2: Quality measures for the photometric redshift determination as measured on the test sample.
Figure 6: Probability densities generated by the CGAN proposed on this paper compared with the true spectroscopic redshift measured at the test sample.
Both CGAN and Random Forest Regressor show a similar level of performance. As a comparison with other quality tests from other photometric redshift algorithms made with DES data, we see that the fraction of outliers of the proposed CGAN is at a comparable level with the 15 photometric redshift algorithms proposed at Table 8 from Sánchez et al. (2014). Nevertheless, our proposed CGAN has one order of magnitude better precission for the mean and median bias, reaching only 3 out of the 15 algorithms a comparable magnitude to the one obtained by the proposed CGAN of this paper: TPZ, SkyNet and ANNz.
As an additional by-product, we can inspect the redshift probability density inferred by the CGAN for individual galaxies. At Figure 6, we can see 9 randomly selected galaxies, comparing the inferred probability density function with the respective true spectroscopic redshift. Testing if the probability densities of the redshift inferred for individual galaxies can not be directly sampling the spectroscopic redshift, as we only have one single galaxy. Nevertheless, we can make use of the Chebyshev inequality to probe that the photometric redshift produced by our CGAN is indeed a probability distribution. Chebyshev inequality states that if $P$ is any probability density with mean photometric redshift $\mu$ and variance $\sigma$, then, $\forall a>0$ the following inequality holds (Canavos et al., 1988),
$P(|\hat{y}-\mu|\geq a\sigma)\leq\frac{1}{a^{2}}.$ (8)
Thus, for each individual galaxy, its photometric redshift for the probability density is generated. From this probability density, the mean and the standard deviation is computed. Nine different values of $a$ are considered beforehand and the number of galaxies for which the inequality $|\hat{y}-\mu|\geq a\sigma$ holds is computed. By Chebyshev inequality, the fraction of outliers for every $a$ shall be lower or equal to $1/a^{2}$.
At Figure 7, we can see the fraction of outliers measured by the CGAN compared with the upper bound imposed by the Chebyshev inequality. We can see that the fraction of outliers is below the Chebyshev bound for all values of $a$. Thus, we can conclude that the photometric redshift probability densities inferred by the CGAN, are indeed a probability density. Neverhteless, we can not demonstrate how accurate these probability densities represent the real spectroscopic redshift for individual galaxies.
Figure 7: Validation of the Chebyshev inequality for the test data sample.
4 Conclusions
On this paper, it was presented a new machine-learning technique for photometric redshift estimation using Conditional Generative Adversarial Networks (CGANs), was presented. Propsed CGAN techique was tested using Dark Energy Survey Y1 data and compared with the classic machine-learning technique the Random Forest Regressor.
Results obtained at CGAN approach, exhibit a similar performance for determining the mean photometric redshift as the Random Forest Classifier. Whereas Random Forest Regressor can only provide a the expected redshift for each individual galaxy, the proposed CGAN approach is able to provide a full probability density function. This additional capability of the CGAN, allows for a better control for the impact of thick tails of the distribution, making this type of photometric redshift algorithmic approach more suitable for cosmological probes that are specially sensitive to thick tails at photometric redshifts -such as number-counts magnification-.
In addition to the value added, CGAN demonstrates a superior computing performance over Random Forest Regressor, as modern Deep Learning frameworks -such as PyTorch-, have a strong parallelization optimization for the determination of the weights of the tensors defining the neural networks, whereas algorithms such as the Random Forest Regressor are not parallelization. This, makes the CGAN approach superior from the computing performance point of view, which is specially suitable for large Cosmology surveys.
It is worth remarking that the results of this work do not constitute a full out-of-the-box software for photometric redshift determination, rather than a proof of concept and a feasibility analysis. This algorithm as still room for improvement in terms of calibration sample selection and the inclusion of additional features.
Accuracy metrics obtained by CGAN, exhibit a perfomance ranking with the best photometric redshift trained on similar calibration data. Performance and accuracy metrics obtained by the CGAN approach described here, is strongly limited by the resolution of the band-pass filters used for the photometry and the galaxy sample used. On this dataset, only four band-pass filters $griz$ are available, thus, much better photometric redshift could be obtained by using a higher resolution apparatus, like the 40-band-pass filter PAUCam.
Future work will explore the possibility of including a finer selection by population of galaxies (main-sequence, LRGs, etc.) that can be included as an additional parameter of the CGAN. Additional features to be explored will also include an evaluation of the inclusion of additional photometric features of the galaxies, such as shape parameters.
Acknowledgements
Author wants to thank to Departamento de Computación y Tecnología of Universidad Europea de Madrid for providing access to computing resources at the LORCA Cluster.

### v2

@@
 Abstract
-Accurate and reliable photometric redshifts determination is one of the key aspects for wide-field photometric surveys.
-Determination of photometric redshift for galaxies, has been traditionally solved by use of machine-learning and artificial intelligence techniques trained on a calibration sample of galaxies, where both photometry and spectrometry are determined.
+Accurate and reliable photometric redshift determination is one of the key aspects for wide-field photometric surveys.
+Determination of photometric redshift for galaxies, has been traditionally solved by use of machine-learning and artificial intelligence techniques trained on a calibration sample of galaxies, where both photometry and spectrometry are available.
 On this paper, we present a new algorithmic approach for determining photometric redshifts of galaxies using Conditional Generative Adversarial Networks (CGANs).
-Proposed CGAN implementation, approaches photometric redshift determination as a probabilistic regression, where instead of determining a single value for the estimated redshift of the galaxy, a full probability density is computed.
-The methodology proposed, is tested with data from Dark Energy Survey (DES) Y1 data and compared with other existing algorithm such as a Random Forest regressor.
+The proposed implementation is able to determine both point-estimation and probability-density estimations for photometric redshifts.
+The methodology is tested with data from Dark Energy Survey (DES) Y1 data and compared with other existing algorithm such as a Mixture Density Network (MDN).
+Although results obtained show a superiority of MDN, CGAN quality-metrics are close to the MDN results, opening the door to the use of CGAN at photometric redshift estimation.
 keywords:
 Conditional Generative Adversarial Networks , photometric-redshift , galaxy-surveys
-journal: Astronomy and Computing
+journal: New Astronomy
 […]
@@
 At photometric surveys, the redshift of galaxy spectra is inferred by measuring the brightness of galaxies at broad-band filters instead of determining the Doppler shift of their spectra with a high-resolution spectrometer.
-The usual approach for translating from the brightness measured at the broad-band filters to a redshift that is the closest as possible to the potential actual spectroscopic redshift, the use of machine-learning and artificial intelligence techniques are used.
+The usual approach for translating from the brightness measured at the broad-band filters to a redshift, is the use of machine-learning and artificial intelligence.
 These techniques, make use of a calibration sample of galaxies with known both the photometry and the high-resolution spectra.
-This calibration sample, is used by the machine-learning algorithms as a source for indenfying and discovering patterns relating the brightness at the different broad-band filters with the spectroscopic redshift.
-Previous machine-learning algorithms for photometric redshift determination have included: template-fitting (Schaap and van de Weygaert, 2000; Bolzonella et al., 2011; Benítez, 2011; Arnouts and Ilbert, 2011; Laur et al., 2022), neural networks (Collister and Lahav, 2004; Sadeh et al., 2016; Mahmud Pathi et al., 2024; Hoyle, 2016; Chunduri and Mahesh, 2023; Carrasco Kind and Brunner, 2013; Almosallam et al., 2016; Cavuoti et al., 2017), boosted decision trees (Gerdes et al., 2010)-, convolutional neural networks (D’Isanto and Polsterer, 2018; Schuldt et al., 2021), bayesian neural networks (Lima et al., 2022), random forest (Lu et al., 2023), recurrent neural networs (Luo et al., 2024) and nearest neighbours (Graham et al., 2018).
+This calibration sample, is used by the machine-learning algorithms as a source for indentifying and discovering patterns relating the brightness at the different broad-band filters with the spectroscopic redshift.
+Previous machine-learning algorithms for photometric redshift determination have included: neural networks (Collister and Lahav, 2004; Sadeh et al., 2016; Mahmud Pathi et al., 2024; Hoyle, 2016; Chunduri and Mahesh, 2023; Carrasco Kind and Brunner, 2013; Almosallam et al., 2016; Cavuoti et al., 2017), boosted decision trees (Gerdes et al., 2010), convolutional neural networks (D’Isanto and Polsterer, 2018; Schuldt et al., 2021), bayesian neural networks (Lima et al., 2022), random forest (Lu et al., 2023), recurrent neural networs (Luo et al., 2024) and nearest neighbours (Graham et al., 2018).
 A good systematic review on the different types of photometric redshift algorithms can be found at Sánchez et al.
@@
 (2019) and Newman and Gruen (2022).
-As it can be identified at literature review, the most widely used machine-learning technique at photometric redshift estimation are neural networks at their many flavours (dense, recurrent and convolutional), with a strong presence of feed-forward neural networks composed by dense layers.
-Feed-forward neural networks composed by dense layers are a great computational technique for solving the photometric redshift estimation task, as they constitute an excellent regression technique for multivariate analysis that -if provided a proper loss function-, can minimize the errors between the inferred redshift and the actual spectroscopic redshift of the calibration sample without the need for providing an explicit functional dependence between, template or prior assumption (Bishop, 2006; Goodfellow et al., 2016).
-Nevertheless, the estimation of these photometric redshifts constitute a point estimate of the spectroscopic redshift instead of a probability density function.
-A good point estimate is necessary for some cosmological probes -such as Baryon Acoustic Oscillation (Crocce et al., 2018)-, but for other cosmological probes require to have a good control of the tails of the probability distribution function -such as number-counts magnification (Garcia-Fernandez et al., 2018)-.
+Previous photometric redshift machine-learning algorithms that are capable of producing a probability density estimation for the redshift, instead of a point estimation, include: classification algorithms (Rau et al., 2015), hierarchical-models (Leistedt et al., 2019) and mixture density networks (MDNs) (Ansari, Zoe et al., 2021; Teixeira et al., 2024; D’Isanto, A.
+and Polsterer, K.
+L., 2018).
+Nevertheless, these type of algorithms suffer from the pathological defect that the probability-density produced directly by these algorithms, can not be directly interpreted as the redshift probability-density and need additional re-calibrations (Dey et al., 2021).
 A new class of neural network that has still not been explored at photometric redshift estimation are the Generative Adversarial Networks (GANs).
@@
 On this paper, we propose the use of CGANs for estimating photometric redshifts using the magnitudes measured at broad-band filters.
-Proposed algorithm is tested with data from the Dark Energy Survey Y1 data that is overlapping with SDSS Stripe-82.
-Results obtained by proposed CGAN are compared with a classic machine-learning technique such as a Random Forest regressor.
+This algorithm is tested with data from the Dark Energy Survey Y1 data that is overlapping with SDSS Stripe-82.
+Results obtained by proposed CGAN are compared with a Mixture Density Network (MDN).
 2 Methodology
+2.1 Conditional Generative Adversarial Network
 Let ${\bf x}_{i}$ be a sample of photometric data from magnitudes measured at broad-band pass filters for the $i$-th galaxy of some data and $y_{i}$ its corresponding known spectroscopic redshift, such that $\{({\bf x}_{i},y_{i})\}_{i=1}^{N_{train}}$ constitutes the data training set.
 Let ${\bf z}_{i}$ be a vector of randomly-generated numbers [^7] Beware that Computer Science standard notation at literature on GANs, name the random vectors with $z$.
-Do not confuse this $z$ with the redshift of the galaxies (that will be either denoted in this paper as $y$ or $\hat{y}$ for spectroscopic or photometric redshifts respectively., associated to the $i-th$ galaxy.
+Do not confuse this $z$ with the redshift of the galaxies (that will be either denoted in this paper as $y$ or $\hat{y}$ for spectroscopic or photometric redshifts respectively., associated to the $i$th galaxy.
 The generator network $G$ constitutes a function such that $\hat{y}_{i}=G({\bf z}_{i}|{\bf x}_{i};\theta_{G})$, where $\hat{y}_{i}$ is the estimated photometric redshift for the set of magnitudes ${\bf x}_{i}$ and $\theta_{G}$ are the weights defining the generator neural network.
-On the other side, the discriminator network $D$ provides a function such that $d_{i}=D(\hat{y}_{i};\theta_{D})$with $d_{i}\in[0,1]$- is a classifier that identifies if $\hat{y}_{i}$ is a real data from the training sample or a synthetic-generated data produced by the generator network and $\theta_{D}$ are the weights defining the discriminator network.
+On the other side, the discriminator network $D$ provides a function such that $p_{i}=D(\hat{y}_{i};\theta_{D})$with $p_{i}\in[0,1]$- is a classifier that identifies if $\hat{y}_{i}$ is a real data from the training sample or a synthetic-generated data produced by the generator network and $\theta_{D}$ are the weights defining the discriminator network.
 The process of training a GAN network constitutes a min-max problem such that $\min_{\theta_{G}}\max_{\theta_{D}}V(D,G)$.
@@
 The functions $g_{f}$ and $f^{*}$ can be chosen freely, provided they are derived from any $f$-divergence (Csiszár and Shields, 2004; Liese and Vajda, 2006; Nguyen et al., 2007; Reid and Williamson, 2011).
-Conversely, $\mathbb{E}_{{\bf x}\sim p_{d}({\bf x})}$, $\mathbb{E}_{y\sim p_{d}(y)}$ and $\mathbb{E}_{{\bf z}\sim p_{d}({\bf z})}$ denote the expected values over $\bf x$, $y$ and $\bf z$ respectively.
+On the other side, $\mathbb{E}_{{\bf x}\sim p_{d}({\bf x})}$, $\mathbb{E}_{y\sim p_{d}(y)}$ and $\mathbb{E}_{{\bf z}\sim p_{d}({\bf z})}$ denote the expected values over $\bf x$, $y$ and $\bf z$ respectively.
 Taking this into account, the loss function to be minimized for the discriminator network is given by
@@
 $\hat{y}_{i}({\bf z}_{j})=G({\bf z}_{j}|{\bf x}_{i}).$ (7)
-If $N_{r}$ different random vectors are considered, the set of values $\{\hat{y}_{i}({\bf z}_{j})\}_{j=1}^{N_{r}}$, will be a random variable distributed according to the underlying conditional probability density of the spectroscopic redshift for the value of magnitudes ${\bf x}_{i}$.
-Thus, provided that $N_{r}$ is large, the distribution of $\hat{y}_{i}({\bf z_{j}})$ is an estimator of the probability density function of the true spectroscopic redshift for that single $i$-th galaxy.
 Proposed topology of the neural network for the generator is a sequence of 3 fully connected layers, where the first two layers are followed by a Batch Normalization layer and a ReLU activation function.
@@
 The output of the generator network is directly the inferred photometric redshift $\hat{y}$.
-For the discriminator neural network a sequence of 3 fully connected layers, where the first two layers are followed by a Batch Normalization layer and a ReLU activation function.
+The discriminator neural network, is a sequence of 3 fully connected layers, where the first two layers are followed by a Batch Normalization layer and a ReLU activation function.
 Input neurons of the first fully connected layer has $5$ neurons with $D_{DIM}$ output neurons.
@@
 The output of the discriminator network is a decimal number between 0 and 1 that indicates the probability of the input galaxy to be real.
-Topology of generator and discriminator neural networks can be seen at Figure 1.
-Figure 1: Topology of the neural networks used. Blue boxes denote the fully connected layers, green boxes denote the normalization layers and red boxes denote the activation functions.
-In order to compare the results of the proposed CGAN with classic machine-learning approaches, a Random Forest regressor is also employed (Breiman, 2001)
+Topology of generator and discriminator neural networks can be seen at Figure 1 and a diagram illustrating how the Generator and Discriminator networks interact, can be seen at Figure 2.
+Figure 1: Topology of the Generator and Discriminator networks of the CGAN model.
+Gray neurons $\bf x_{i}$, denote the input feature vector containing the four MAG_AUTO magnitudes.
+Blue neurons indicate the $Z_{DIM}$ neurons relative to the input random vector $\bf z_{i}$.
+Green neuron $\hat{y}_{i}$ indicate the predicted photometric redshift, which is both the output of the Generator and part of the input of the Discriminator.
+The orange neuron $p_{i}$ is the probability of some observation to be synthetic data generated by the Generator or data from the real sample.
+Purple neurons are the $G_{DIM}$ and $D_{DIM}$ of the hidden layers.
+Figure 2: Interaction of the Generator and Discriminator networks of the CGAN approach.
+A tensor of random noise $z_{i}$ (gray box) is fed into the Generator network, producing an estimation of the photometric redshift.
+Tensors containing real spectroscopic redshift ($y_{i}$) and the estimated photometric redshifts ($\hat{y}_{i}$) are fed into the Discriminator network, which tries to guess which records are real spectroscopic data and which records are produced by the Generator network given the magnitude data $x_{i}$ (blue box).
+More recent works on GANs, have made extensively use of Wasserstein Arjovsky et al.
+(2017) formalism (WGANs) instead of $f$-divergences, showing superior empirical performance and easier convergence of GANs (Brock et al., 2019; Karras et al., 2018).
+Nevertheless, WGANs prevent the interpretation of the outputs as a probability density (Song and Ermon, 2020).
+Thus, Wasserstein formalism has been discarded for this work.
+2.2 Mixture Density Network
+Mixture Density Networks (MDNs), are a type of machine-learning model that combines a neural-network and a parametric mixture model (Bishop, 1994; McLachlan et al., 2019; Yu et al., 2012).
+Thus, instead of producing a point estimation, MDNs learn the conditional probability as a linear combination of a finite set of individual probability distributions.
+To do so, the neural-network of the MDN aims the prediction of the parameters that characterizes the probability distributions and the mixing coefficients.
+The most common probability density used, is a Gaussian.
+Thus, the neural network aims to determine the set of parameters $\{(\mu_{i},\sigma_{i},\pi_{i})\}_{i=0}^{N_{g}}$, where $\mu_{i}$ is the mean of the Gaussian, $\sigma_{i}$ the standard-deviation and $\pi_{i}$ the mixing coefficients (such that $\sum_{i=0}^{N_{g}}\pi_{i}=1$).
+On the other side, $N_{g}$ is the number of gaussians on the mix and shall be a configuration parameter of the model that must be fixed.
+In order to compare with a Mixture Density Network, in this work, we will be using the MDN proposal from Ansari, Zoe et al.
+(2021), where we took the code from companion GitHub [^8] https://github.com/ZoeAnsari/MixtureModelsForPhotometricRedshifts and ported it from Keras to PyTorch.
+Configuration parameters of the MDN model are kept the same as the original formulation, except the number of neurons of the input layer -where in this work 4 neurons are used instead of the original 22 input neurons-.
+This is made in order to adapt original Ansari, Zoe et al.
+(2021) code to the 4 magnitude features of this dataset, instead of the 22 magnitude features used originally at the paper.
+Thus, MDN model is composed by 30 mixture Gaussians, one input fully-connected layer of 4 neurons and a hidden fully-connected layer of 22 neurons.
 3 Data Analysis
@@
 Data was taken from the public data releases DR1 (Abbott et al., 2018; Drlica-Wagner et al., 2018) from The Dark Energy Survey Collaboration that are available at the NCSA repository [^9] https://des.ncsa.illinois.edu/releases/y1a1.
-Code for the CGAN was implemented under Python [^9] https://www.python.org using PyTorch [^10] https://pytorch.org (Paszke et al., 2019), whereas code for the Random Forest Regressor was implemented using Scikit-Learn [^11] https://scikit-learn.org (Pedregosa et al., 2011).
+Code for the CGAN was implemented under Python [^10] https://www.python.org using PyTorch [^11] https://pytorch.org (Paszke et al., 2019).
 Full code of this analysis can be found at author’s GitHub repository [^12] https://github.com/mgarciafernandez-uem/CGAN-photoz.
@@
 Hyperparameters of the model defining the sizes of the dense layers of generator ($G_{DIM}$) and the discriminator ($D_{DIM}$) along with the size the random vector ($Z_{DIM}$) can be seen at Table 1.
-Training strategy used the Adam optimizer with a learning rate parameter of $lr=10^{-4}$ for both generator and discriminator network using $10000$ training epochs, using a step learning-rate approach reducing the learning rate by a factor of $0.2$ every $2000$ training epochs.
+Training strategy used the Adam optimizer with an initial learning rate parameter of $lr=10^{-4}$ for both generator and discriminator network using $10000$ training epochs, using a step learning-rate approach reducing the learning rate by a factor of $0.2$ every $2000$ training epochs.
 Training strategy is the same for the generator and the discriminator networks.
@@
 A plot showing the Mean Square Error (MSE) between the inferred photometric redshift and the true spectroscopic redshift is also displayed.
-A comparison of the observed distribution of inferred redshift of galaxies for the test sample is compared with the true spectroscopic redshift can be seen at Figure 4.
-In addition, a comparison of the dispersion between the true spectroscopic redshift and the inferred photometric redshift can be see at Figure 5.
-From these plots, we can see that the photometric redshifts inferred by the proposed CGAN approach can accurately trace the true spectroscopic redshift distribution of the test sample of galaxies, reaching an agreement similar to the classic Random Forest Regressor.
-In addition, we can see that there is a similar level of dispersion on the photometric redshifts inferred by the CGAN approach and the Random Forest Regressor.
-It is important to remark that no significant biases or asymmetries are spotted.
-We see that dispersion of inferred photometric redshift randomly fluctuate around the true spectroscopic redshift for all the range of values of redshift.
-Figure 4: Distribution of the photometric redshift of galaxies inferred from both proposed CGAN approach and the classic Random Forest algorithm.
-Both distributions are compared with the true spectroscopic redshift (black solid-line).
-Measurements are made on the test sample of galaxies.
-Figure 5: Dispersion between the photometric redshift inferred by proposed CGAN and classic Random Forest algorithm and the true spectroscopic redshift.
-The black-dashed line displays those redshifts where the photometric redshift is equal to the spectroscopic redshift.
-Measurements are made on the test sample of galaxies.
-Quality metrics for photometric redshift determination are based on the ones proposed at Sánchez et al.
-(2014) and can be found at Table 2.
-Used quality metrics for photometric comparison, are: the mean bias between the photometric redshift and the spectroscopic redshift ($\bar{\Delta z}$), the median of the bias ($\Delta z_{50}$), the standard deviation of the bias ($\sigma_{\Delta z}$), the fraction of galaxies with spectroscopic redshift outside of the two-standard deviation range ($out_{2\sigma}$) and the fraction of galaxies with spectroscopic redshift outside of the three-standard deviation range ($out_{3\sigma}$).
-Parameter Description CGAN Random Forest
-$\bar{\Delta z}$ Mean bias -0.0046 -0.0030
-$\Delta z_{50}$ Median of the bias -0.0046 -0.0055
-$\sigma_{\Delta z}$ Standard deviation of the bias 0.092 0.056
-$out_{2\sigma}$ Outliers beyond 2 standard deviations 4% 5%
-$out_{3\sigma}$ Outliers beyond 3 standard deviations 2% 2%
-Table 2: Quality measures for the photometric redshift determination as measured on the test sample.
-Figure 6: Probability densities generated by the CGAN proposed on this paper compared with the true spectroscopic redshift measured at the test sample.
-Both CGAN and Random Forest Regressor show a similar level of performance.
-As a comparison with other quality tests from other photometric redshift algorithms made with DES data, we see that the fraction of outliers of the proposed CGAN is at a comparable level with the 15 photometric redshift algorithms proposed at Table 8 from Sánchez et al.
+MDN training process involves 5000 training epochs with an initial learning-rate of $lr=10^{-4}$. Losses and mean-square-error (MSE) for the MDN can be seen at Figure 5
+Figure 5: Losses and mean-square-error (MSE), for the Mixture Density Network (MDN).
+A comparison of the dispersion between the true spectroscopic redshift and the inferred photometric redshift can be see at Figure 6.
+From these plots, we can see that the photometric redshifts inferred by the proposed CGAN approach and the MDN can accurately trace the true spectroscopic redshift distribution of the test sample of galaxies.
+Figure 6: Comparison of the photometric redshift of galaxies and the spectroscopic redshift.
+Distributions are displayed for proposed Conditional Generative Adversarial Network (CGAN) approach and the Mixture Density Network (MDN).
+White-dashed line is an eye-guide for the identity.
+For determining the quality metric both for point-estimation and probability density estimation, we use the approach from Teixeira et al.
+(2024), where these metrics are all determined over the test sample of galaxies which have not been seen neither by the CGAN neither the MDN.
+3.1 Point-estimation quality metrics
+Point-estimation quality metrics involve: the mean absolute bias ($\bar{|\Delta z|}$), the Normalized Median Absolute Deviation ($\sigma_{NMAD}$) and the outliers ratio $\eta$.
+Metrics are computed in 20 spectroscopic redshift at the interval $0\leq z_{sp}\leq 0.8$.
+The Normalized Median Absolute Deviation is given by
+$\sigma_{NMAD}=1.48\times median\left(\frac{\Delta z-median(\Delta z)}{1+z}\right),$ (8)
+whereas the outlier ratio ($\eta$) is defined as the ratio of galaxies with
+$\left|\frac{\Delta z}{1+z}\right|>0.15.$ (9)
+The confidence intervals for these metrics are computed at the $95\%$ confidence level by using bootstrapping.
+Bootstrapping is implemented by building 1000 samples of the test dataset with each sub-sample having the same number of galaxies as in the full test dataset, where each sample is generated by picking randomly with replacement galaxies from the test dataset.
+Thus, for each subsample at the bootstrap, each point-estimation quality-metric is computed.
+Then, this set of 1000 metrics is used to determine the respective $2.5\%$ and $97.5\%$ quantiles, which stand for the lower and upper edges of the confidence intervals.
+Graphics for these quality metrics can be seen at Figure 7, from which we can see that both CGAN and MDN models have a close performance at quality metrics, being the MDN more accurate than the CGAN, being the latter more prone to outliers.
+Figure 7: Point-estimation quality-metrics comparison for CGAN and MDN.
+Left, shows the mean absolute bias $\bar{|\Delta z|}$.
+Center shows the $\sigma_{NMAD}$.
+Right shows the ratio of outliers $\eta$.
+Solid line stands for the mean value per redshift bin.
+Shaded area stands for the 95% confidence interval for that redshift bin.
+Confidence intervals are computed using bootstrap.
+3.2 Probability-density-function quality metrics
+Given a galaxy with spectroscopic redshift $z_{sp}$, photometric redshift $z_{ph}$ with associated PDF $\phi(z)$, its PIT is defined as (Dawid, 1984; Lima et al., 2022; Polsterer et al., 2016)
+[…]
+As stated by Mucesh et al. (2021), a properly calibrated PDF, will produce an uniform distribution of PIT values for a large sample of galaxies.
+The Odds (Lima et al., 2022) for a galaxy with PDF $\phi(z)$, with photometric redshift $z_{ph}$ is defined as
+[…]
+where $\xi\in\mathbb{R}$ is a constant parameter that must be fixed before-hand, that is fixed to $\xi=0.06$ taking the same value as Teixeira et al.
 (2024).
-Nevertheless, our proposed CGAN has one order of magnitude better precission for the mean and median bias, reaching only 3 out of the 15 algorithms a comparable magnitude to the one obtained by the proposed CGAN of this paper: TPZ, SkyNet and ANNz.
-As an additional by-product, we can inspect the redshift probability density inferred by the CGAN for individual galaxies.
-At Figure 6, we can see 9 randomly selected galaxies, comparing the inferred probability density function with the respective true spectroscopic redshift.
-Testing if the probability densities of the redshift inferred for individual galaxies can not be directly sampling the spectroscopic redshift, as we only have one single galaxy.
-Nevertheless, we can make use of the Chebyshev inequality to probe that the photometric redshift produced by our CGAN is indeed a probability distribution.
-Chebyshev inequality states that if $P$ is any probability density with mean photometric redshift $\mu$ and variance $\sigma$, then, $\forall a>0$ the following inequality holds (Canavos et al., 1988),
-$P(|\hat{y}-\mu|\geq a\sigma)\leq\frac{1}{a^{2}}.$ (8)
-Thus, for each individual galaxy, its photometric redshift for the probability density is generated.
-From this probability density, the mean and the standard deviation is computed.
-Nine different values of $a$ are considered beforehand and the number of galaxies for which the inequality $|\hat{y}-\mu|\geq a\sigma$ holds is computed.
-By Chebyshev inequality, the fraction of outliers for every $a$ shall be lower or equal to $1/a^{2}$.
-At Figure 7, we can see the fraction of outliers measured by the CGAN compared with the upper bound imposed by the Chebyshev inequality.
-We can see that the fraction of outliers is below the Chebyshev bound for all values of $a$.
-Thus, we can conclude that the photometric redshift probability densities inferred by the CGAN, are indeed a probability density.
-Neverhteless, we can not demonstrate how accurate these probability densities represent the real spectroscopic redshift for individual galaxies.
-Figure 7: Validation of the Chebyshev inequality for the test data sample.
+A distribution of Odds peaked towards large values, indicates that the PDF produced are narrow and the inferred photometric redshift is around the most probable value, indicating a reliable PDF.
+On the other side, low Odds value, indicates that PDFs are broad.
+Coverage-Test (Dalmasso et al., 2020; Hermans et al., 2020), is computed by taking for each galaxy its spectroscopic redshift ($z_{sp}$) and the PDF ($\phi(z)$).
+For a prefixed value of confidence level $\alpha-1$, the edges $z_{l},z_{u}$ of the symmetric interval enclosing a probability of $1-\alpha$ is computed by using
+[…]
+Thus, the ratio of galaxies with spectroscopic redshift $z_{sp}$ at the interval such that $z_{l}\leq z_{sp}\leq z_{u}$, must be exactly equal to $1-\alpha$.
+A ratio of galaxies within the interval lower than expected, indicates that the PDFs produced are narrower than expected pointing an overconfidence of the algorithm on the determination of the PDFs for the galaxies.
+On the other side, higher ratios of galaxies than expected, indicate a broader PDFs than expected, pointing a underconfidence of on the determination of the PDFs.
+Graphics for these quality-metrics can be found at Figure 8.
+We can see that both the PIT and Odds distributions of the CGAN and the MDN approaches are very similar.
+On the other side, we can see on the Coverage-Test, that the measured confidence values for the CGAN, show that both probability densities are skewed towards overconfidence.
+Figure 8: Probability density quality metrics comparison for CGAN and MDN. Left shows the PIT. Center shows the Odds. Right shows the credibility-diagram.
+In addition, in order to test the capacity of the algorithm to recover the probability density function, the distribution of the spectroscopic redshift of the galaxies is compared with the distribution computed by stacking all the probability densities of the individual galaxies produced by both algorithms.
+The result can be seen at Figure 9, where we can see that the stacking of the probability densities show a similar description of the underlying data distribution, with the MDN being more closer to spectroscopic data than the CGAN.
+Figure 9: Comparison of the probability density of the spectroscopic redshift with the probability density inferred from the stacking of the probability densities of individual galaxies for CGAN and MDN algorithms.
 4 Conclusions
-On this paper, it was presented a new machine-learning technique for photometric redshift estimation using Conditional Generative Adversarial Networks (CGANs), was presented.
-Propsed CGAN techique was tested using Dark Energy Survey Y1 data and compared with the classic machine-learning technique the Random Forest Regressor.
-Results obtained at CGAN approach, exhibit a similar performance for determining the mean photometric redshift as the Random Forest Classifier.
-Whereas Random Forest Regressor can only provide a the expected redshift for each individual galaxy, the proposed CGAN approach is able to provide a full probability density function.
-This additional capability of the CGAN, allows for a better control for the impact of thick tails of the distribution, making this type of photometric redshift algorithmic approach more suitable for cosmological probes that are specially sensitive to thick tails at photometric redshifts -such as number-counts magnification-.
-In addition to the value added, CGAN demonstrates a superior computing performance over Random Forest Regressor, as modern Deep Learning frameworks -such as PyTorch-, have a strong parallelization optimization for the determination of the weights of the tensors defining the neural networks, whereas algorithms such as the Random Forest Regressor are not parallelization.
-This, makes the CGAN approach superior from the computing performance point of view, which is specially suitable for large Cosmology surveys.
-It is worth remarking that the results of this work do not constitute a full out-of-the-box software for photometric redshift determination, rather than a proof of concept and a feasibility analysis.
-This algorithm as still room for improvement in terms of calibration sample selection and the inclusion of additional features.
-Accuracy metrics obtained by CGAN, exhibit a perfomance ranking with the best photometric redshift trained on similar calibration data.
-Performance and accuracy metrics obtained by the CGAN approach described here, is strongly limited by the resolution of the band-pass filters used for the photometry and the galaxy sample used.
-On this dataset, only four band-pass filters $griz$ are available, thus, much better photometric redshift could be obtained by using a higher resolution apparatus, like the 40-band-pass filter PAUCam.
-Future work will explore the possibility of including a finer selection by population of galaxies (main-sequence, LRGs, etc.) that can be included as an additional parameter of the CGAN.
-Additional features to be explored will also include an evaluation of the inclusion of additional photometric features of the galaxies, such as shape parameters.
-Acknowledgements
+On this paper, it was presented a new machine-learning technique for photometric redshift estimation using a Conditional Generative Adversarial Network (CGAN).
+Propsed CGAN techique was tested using Dark Energy Survey Y1 data and compared with a Mixture Density Network (MDN).
+Both point-estimation and probability-density quality-metrics indicate CGAN performance is close to the MDN algorithm.
+Although these metrics clearly demonstrate that MDN performs better than the proposed CGAN, this work constitutes a proof of concept indicating that the use of CGAN with the $f$-divergence formalism is an algorithm that can be used for photometric redshift determination, opening the door to further developments on the use of CGANs at this field.
+Future work will explore the possibility of including a finer selection by population of galaxies (main-sequence, LRGs, etc.) that can be included as an additional parameter of the CGAN that shall be computed as a previous step to the photometric redshift CGAN.
+Acknowledgments
 Author wants to thank to Departamento de Computación y Tecnología of Universidad Europea de Madrid for providing access to computing resources at the LORCA Cluster.